# Exploring Dark Forces with Multimessenger Studies of Extreme Mass Ratio Inspirals

### v1

@@ 1 Introduction
This widely used framework has historically served to parametrize deviations from GR and the inverse square law [1, 2, 3], the effects of light string moduli [4, 5, 6, 7] and, in more recent times, as a template to study light mediators in dark sector models [8, 9, 10, 11, 12]. Interpreted purely in terms of deviations from GR (i.e., without reference to dark sector physics), this formalism describes, for example, models of extra dimensions (for small values of $\lambda\ll 10^{-3}$m) [13, 14] and the linearized regime of massive gravity theories (for large values of $\lambda$) [1]. In the study of dark sectors, on the other hand, $\tilde{\alpha}^{\prime}$ takes on the role of the dark sector coupling and $m_{V}$ the role of the dark mediator mass, while the Yukawa interaction is assumed to be sourced by the accumulation of dark charge on the central and secondary objects. Indeed, a host of interaction potentials (not just the Yukawa potential) are possible, depending on the nature of the mediator being exchanged (we refer to [15] for a unified treatment of such exchange interactions, and [16, 17] for recent work along these lines). We will mainly restrict our attention to long-ranged Yukawa potentials in this work.
@@
Compact object mergers in particular present unique opportunities: the population distributions and the GW waveforms emanating from these systems are sensitive to modifications of the type parametrized in Eq. (1.1). Different classes of compact binary systems enable different scales $\lambda$ to be probed, depending on the characteristic frequency of the GW emission. The dark sector interpretation of Eq. (1.1), which is our main interest, works under the assumption that both the interacting objects accumulate dark charge. In the case of inspiraling neutron stars, such accumulation could result from dark matter capture; in the case of solar or super massive black holes, such accumulation could result from spike formation, or from the fact that such black holes are charged under millicharged dark fermions [23]. Here, too, the recent literature is substantial, and we refer to [8, 9, 10, 11] for some representative studies.
@@ 3 Results
We note that several prior studies of dark sectors with binary inspirals have concentrated on constraining either the strength of the dipole radiation term, or the charge of the remnant. For example, [57, 23, 58, 11] consider black hole binary inspirals with massless mediators and constrain the dipole radiation term from the early pre-merger inspiral stage. The charge of the merger remnant can be constrained from the detection of the quasinormal modes. Dipole radiation in the context of EMRIs has been considered by [59, 60, 61]. Our focus, as stated before, will instead be on constraining the dark force in the limit of vanishing dipole radiation, relying instead on independent mass measurements of the SMBH.

### v2

@@ 1 Introduction
This widely used framework has historically served to parametrize deviations from GR and the inverse square law [1, 2, 3], the effects of light string moduli [4, 5, 6, 7] and, in more recent times, as a template to study light mediators in dark sector models [8, 9, 10, 11, 12, 13]. Interpreted purely in terms of deviations from GR (i.e., without reference to dark sector physics), this formalism describes, for example, models of large extra dimensions (for small values of $\lambda\ll 10^{-3}$m) [14, 15] as well as braneworld [16, 17, 18], and the linearized regime of massive gravity theories (for large values of $\lambda$) [1]. In the study of dark sectors, on the other hand, $\tilde{\alpha}^{\prime}$ takes on the role of the dark sector coupling and $m_{V}$ the role of the dark mediator mass, while the Yukawa interaction is assumed to be sourced by the accumulation of dark charge on the central and secondary objects. Indeed, a host of interaction potentials (not just the Yukawa potential) are possible, depending on the nature of the mediator being exchanged (we refer to [19, 20] for a unified treatment of such exchange interactions, and [21, 22] for recent work along these lines). We will mainly restrict our attention to long-ranged Yukawa potentials in this work.
@@
Compact object mergers in particular present unique opportunities: the population distributions and the GW waveforms emanating from these systems are sensitive to modifications of the type parametrized in Eq. (1.1). Different classes of compact binary systems enable different scales $\lambda$ to be probed, depending on the characteristic frequency of the GW emission. The dark sector interpretation of Eq. (1.1), which is our main interest, works under the assumption that both the interacting objects accumulate dark charge. In the case of inspiraling neutron stars, such accumulation could result from dark matter capture; in the case of solar or super massive black holes, such accumulation could result from spike formation, or from the fact that such black holes are charged under millicharged dark fermions [28]. Here, too, the recent literature is substantial, and we refer to [8, 9, 10, 11, 12] for some representative studies.
@@ 3 Results
We note that several prior studies of dark sectors with binary inspirals have concentrated on constraining either the strength of the dipole radiation term, or the charge of the remnant. For example, [62, 28, 63, 64, 11] consider black hole binary inspirals with massless mediators and constrain the dipole radiation term from the early pre-merger inspiral stage. The charge of the merger remnant can be constrained from the detection of the quasinormal modes. Dipole radiation in the context of EMRIs has been considered by [65, 66, 67]. Our focus, as stated before, will instead be on constraining the dark force in the limit of vanishing dipole radiation, relying instead on independent mass measurements of the SMBH.
@@ 4 Conclusions
In this study, we have assumed that the corrections from environmental effects are negligible. SMBHs are known to exist in dense astrophysical environments, which may influence the dynamics of EMRIs [76, 77, 78]. Key processes such as gravitational pull from matter, accretion onto EMRI components, dynamical friction, and planetary migration can impact the gravitational waveform, potentially leading to biased parameter estimation and reduced detectability of EMRIs. We direct readers to [79] for information on the maximum level of corrections due to various environmental effects over a typical LISA mission. The environmental corrections are expected to be minimal for SMBHs surrounded by thick disks, but more substantial for those in thin disks, which are predominantly found in active galactic nuclei at high redshifts. Although SMBHs in thin disks contribute up to only a few percent of the EMRIs detectable by LISA with $z\lesssim 1$, potential uncertainties from environmental effects should also be considered when applying our method to these events.